given two angles that form a linear pair, prove that the bisectors of the two angles are perpendicular to one another.
Answer by mananth(16949)

.
Angle POQ +POR = 180 deg. ( Linear Pair)
MO & NO are the angle bisectors
2a+2b=180 deg
2(a+b) = 180 deg
a+b = 90 deg